# Cyclotomic generating functions

### v2

@@ 3.2. Rational products of $q$-integers
(3.2) $f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}=\sum_{k=0}^{\infty}c_{k}q^{k}$
as formal power series in $\mathbb{Z}[[q]]$. Such rational products include the set of basic CGF’s. We examine several properties of such products.
In the next lemma, we state an explicit formula for the coefficients of the expansion of (3.2) generalizing work of Knuth for the number of permutations with $k\leq n$ inversions in $S_{n}$ in [Knu73, p.16]. See also [Sta12, Ex. 1.124] and [OEI23, A008302], and the application to standard Young tableaux in [BKS20b]. Here we use
@@
for all $k\in\mathbb{Z}_{\geq 0}$ and $x\in\mathbb{Z}$, including $x<0$. For all $x\in\mathbb{Z}$, define the empty product $\binom{x}{0}=1$.
Lemma 3.2.
@@
(3.3) $c_{k}=\sum_{|\mu|=k}\prod_{\begin{subarray}{c}i\geq 1\\ m_{i}(\mu)>0\end{subarray}}\binom{M_{i}+m_{i}(\mu)-1}{m_{i}(\mu)}$
where $m_{i}(\mu)$ is the number of parts of $\mu$ of size $i$. Moreover, we may restrict the sum in (3.3) to only those $|\mu|=k$ where for all $i\geq 1$, either $M_{i}>0$ or $m_{i}(\mu)\leq|M_{i}|$.
@@
Equation (3.3) follows using the expansion $(1-q^{i})^{-j}=\sum_{n=0}^{\infty}\binom{j+n-1}{n}q^{in}$, multiplication of ordinary generating functions, and the fact that $\binom{x}{0}=1$. For the restriction, consider a term indexed by $\mu$ with $M_{i}\leq 0$ and $m_{i}(\mu)>|M_{i}|=-M_{i}$. It follows that $M_{i}\leq 0\leq M_{i}+m_{i}(\mu)-1$, so that
$m_{i}(\mu)!\binom{M_{i}+m_{i}(\mu)-1}{m_{i}(\mu)}=(M_{i}+m_{i}(\mu)-1)\cdots M_{i}=0.$
Remark 3.3.
The binomial coefficients $\binom{M_{i}+m_{i}(\mu)-1}{m_{i}(\mu)}$ appearing in (3.3) are positive when $M_{i}>0$. When $M_{i}\leq 0$ and $m_{i}(\mu)\leq|M_{i}|$, the binomial coefficient is non-zero and has sign $(-1)^{m_{i}(\mu)}$. Thus the negative summands in (3.3) are precisely those for which $\mu$ has an odd number of row lengths $i$ such that $M_{i}\leq 0$.
Remark 3.4.
In the particular case when $f(q)$ is a basic cyclotomic generating function, the fact that the expression in (3.3) is non-negative, symmetric, and eventually 00 is quite remarkable. Is there a simplification of the expression in (3.3) that would directly imply non-negativity of the coefficients for any $f(q)\in\Phi^{+}$? Even the case of $[n]_{q}!$ or any of the examples CGF’s mentioned in the Introduction would be of interest.
3.3. Necessary conditions for $q$-integer ratios to be polynomial
@@
(3.4) $f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}$
to yield a polynomial (not necessarily with positive coefficients) is for the multiplicity of each primitive $d$th root of unity to be non-negative. However, both numerator and denominator can be factored uniquely into cyclotomic polynomials, so the polynomiality criterion is equivalent to
(3.5) #$\#\{k:\ell\mid a_{k}\}\geq\#\{k:\ell\mid b_{k}\}\hskip 7.22743pt\qquad\forall\ \ell\in\mathbb{Z}_{\geq 2}.$
Example 3.5.
In Stanley’s $q$-analogue of the hook length formula
@@
the inequality (3.5) reduces to the fact that the $\ell$-quotient of any partition $|\lambda|=n$ has no more than $\lfloor n/\ell\rfloor$ cells. In particular, this gives a quick proof that $[n]_{q}!/\prod_{c\in\lambda}[h_{c}]_{q}\in\mathbb{Z}[q]$, though the stronger result that the coefficients are non-negative is not clear from this approach.
3.4. Necessary conditions for $q$-integer ratio to be a CGF
In addition to satisfying the polynomiality conditions in (3.5) and (3.7), what further restrictions on $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{m}\}$ are required for a rational product of $q$-integers as in (3.4) to yield a cyclotomic generating function? Say $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/\prod_{k=1}^{m}[b_{k}]_{q}=\sum c_{k}q^{k}$. Using the expression for $c_{k}$ in (3.3), $c_{k}\geq 0$ and $\lim_{k\to\infty}c_{k}=0$ are by definition necessary and sufficient to prove $f(q)\in\Phi^{+}$, though these conditions are difficult to use for families of such rational products in practice. We will outline some more direct necessary conditions here. We begin with complete classifications for $m=1,2$.
Lemma 3.6.
@@ 3.4. Necessary conditions for $q$-integer ratio to be a CGF
  2. (2)
    If $f(q)$ is a power series with non-negative coefficients, then $a_{1},a_{2}\in\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{b_{1},b_{2}\}$.
  3. (3)
@@
Conversely, suppose one of (i)-(iv) holds. Then, $f(q)$ is a polynomial if conditions (i) or (ii) hold by 3.6, and (iv) is equivalent to (iii), so it suffices to assume (iii) holds. Use the cyclotomic expansion of $[n]_{q}$ again from (2.3). Since (iii) holds, every cyclotomic divisor of either $[b_{1}]_{q}$ or $[b_{2}]_{q}$ but not both is canceled by a cyclotomic divisor of $[a_{1}]_{q}$, and every divisor of both $[b_{1}]_{q}$ and $[b_{2}]_{q}$ is canceled by divisors of $[a_{1}]_{q}$ and $[a_{2}]_{q}$ together. Thus, (1) holds.
The span condition in statement (2) is proved for all basic CGF’s in 3.9 below. To prove (3), observe that $f(q)\in\Phi^{+}$ implies both the divisibility and span conditions hold by (1) and (2). Conversely, if the any one of the divisibility conditions hold, then $f(q)$ is a polynomial with integer coefficients. If divisibility conditions (i) or (ii) hold, then $f(q)\in\Phi^{+}$ by 3.6, so suppose divisibility condition (iii) and the span condition in (2) holds. Again the case (iv) is similar.
By canceling out common cyclotomic factors, we may suppose without loss of generality that $\gcd(b_{1},b_{2})=1$ and that $a_{1}=b_{1}b_{2}$. Since $a_{2}\in\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{b_{1},b_{2}\}$, there exist positive integers $u,v$ such that $a_{2}=ub_{1}+vb_{2}$. So, we must show that the polynomial $f(q)=[b_{1}b_{2}]_{q}[ub_{1}+vb_{2}]_{q}/[b_{1}]_{q}[b_{2}]_{q}$ has non-negative coefficients.
@@
Let $N(n)=\#\{x,y\in\mathbb{Z}_{\geq 0}:xb_{1}+yb_{2}=n\}$. The coefficient of $q^{n}$ in $f(q)$ is hence
@@
Let $N(n;\alpha,\beta)=\#\{\alpha\leq x,\beta\leq y:xb_{1}+yb_{2}=n\}$. The previous expression becomes
@@
Furthermore, let $N(n;\alpha,\beta;\delta)=\#\{\alpha\leq x<\alpha+\delta,\beta\leq y:xb_{1}+yb_{2}=n\}$. Thus, the expression for the coefficient of $q^{n}$ in $f(q)$ becomes
@@
Assume there exist $(x_{0},y_{0})\in\mathbb{Z}^{2}$ such that $x_{0}b_{1}+y_{0}b_{2}=n$, since otherwise the coefficient of $q^{n}$ in $f(q)$ is 00. Then, all integer solutions of $xb_{1}+yb_{2}=n$ are of the form $x=x_{0}-tb_{2},y=y_{0}+tb_{1}$ for some $t\in\mathbb{Z}$ by the theory of linear Diophantine equations. In particular, there exist unique solutions $(x_{1},y_{1}),(x_{2},y_{2})\in\mathbb{Z}^{2}$ with $0\leq x_{1}<b_{2}$ and $u\leq x_{2}<u+b_{2}$. Hence, $N(n;0,0;b_{2})=\delta_{y_{1}\geq 0}$ and $N(n;u,v;b_{2})=\delta_{y_{2}\geq v}$. Since $x_{1}\leq x_{2}$, we have $y_{1}\geq y_{2}$ as solutions to the linear equation. Thus, $\delta_{y_{2}\geq v}=1\Rightarrow\delta_{y_{1}\geq 0}=1$. Therefore, $N(n;0,0;b_{2})-N(n;u,v;b_{2})$ is non-negative, which completes the proof. ∎
Example 3.8.
For $m=2$, the polynomiality condition alone does not imply non-negative integer coefficients. For example, $[1]_{q}[6]_{q}/[2]_{q}[3]_{q}=\Phi_{6}(q)=q^{2}-q+1$. Here $6\in\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{2,3\}$, but $1$ is not. For $m=3$ and higher, the polynomiality and span conditions do not imply non-negative integer coefficients. For example,
(3.6) $\frac{[4]_{q}[4]_{q}[15]_{q}}{[2]_{q}[3]_{q}[5]_{q}}=q^{13}+q^{11}+q^{10}-q^{9}+2q^{8}+2q^{5}-q^{4}+q^{3}+q^{2}+1.$
However, we do have the following necessary conditions for basic CGF’s related to spans and sums.
Lemma 3.9.
Suppose $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+}$, then the following hold.
@@
    For each $1\leq k\leq m$$a_{k}\in\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{b_{1},\ldots,b_{m}\}$.
@@
    If $a_{1}\leq\cdots\leq a_{m}$ and $b_{1}\leq\cdots\leq b_{m}$, then $a_{1}\geq b_{1}\geq 1$ and $a_{m}\geq b_{m}$.
@@
    Let $\mathcal{X}$ be the random variable corresponding with $f(q)/f(1)$ with mean $\mu=\frac{1}{2}\sum_{k=1}^{m}(a_{k}-b_{k})$ and variance $\sigma^{2}=\frac{1}{6}\sum_{k=1}^{m}(a_{k}^{2}-b_{k}^{2})$. Then, $\log\sigma=\Theta(\log\mu)$. More precisely,
    $\mu/2\leq\sigma^{2}\leq\mu^{2}.$
@@
    From the kurtosis of $\mathcal{X}$, we have
    $\frac{\sum_{i=1}^{m}(a_{i}^{4}-b_{i}^{4})}{\left(\sum_{i=1}^{m}(a_{i}^{2}-b_{i}^{2})\right)^{2}}\leq\frac{5}{3}.$
@@
    (3.7) $\sum_{k=1}^{m}a_{k}^{d}\geq\sum_{k=1}^{m}b_{k}^{d},$
    where the inequality is strict if the polynomial is non-constant. For example, when $d=1$ this inequality is equivalent to noting the degree of $f(q)$ is $\sum_{k=1}^{m}a_{k}-b_{k}\geq 0$.
@@
The support of the power series $\prod_{k=1}^{m}1/(1-q^{b_{k}})$ is $\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{b_{1},\ldots,b_{m}\}$. Since $f(q)$ has non-negative coefficients and constant term $1$, comparing supports gives $\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{a_{1},\ldots,a_{m}\}\subset\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{b_{1},\ldots,b_{m}\}$, which yields (i).
(ii) From (i), it follows that $a_{1}\geq\min\{b_{1},\ldots,b_{m}\}=b_{1}$. From (3.5) at $\ell=b_{m}$, there is some $a_{k}$ such that $b_{m}\mid a_{k}\leq a_{m}$, so $b_{m}\leq a_{m}$.
@@
(iv) Jensen’s Inequality for convex functions gives the inequality of central moments
$\alpha_{2}^{2}=\mathbb{E}[(\mathcal{X}-\mu)^{2}]^{2}\leq\mathbb{E}[(\mathcal{X}-\mu)^{4}]=\alpha_{4}.$
In terms of cumulants, this says $\kappa_{2}^{2}\leq\kappa_{4}+3\kappa_{2}^{2}$, which simplifies to the stated expression by (2.5).
(v) The inequality follows directly from the formula for cumulants of a rational product of $q$-integers (2.5) and the sign constraints in 2.9. ∎
Remark 3.10.
@@
Remark 3.11.
One might wonder if $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+}$ implies $a_{k}\geq b_{k}$ for all $k$ when sorted as in 3.9(ii). However, this is false, though empirically counterexamples are rare. For instance,
@@
is one of only three counterexamples with $5$ $q$-integers in the numerator and denominator and maximum entry $\leq 14$, out of $956719$ cyclotomic generating functions satisfying these conditions. On the other hand, we propose the following.
Conjecture 3.12.
Suppose $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+}$$a_{1}\leq\cdots\leq a_{m}$,  $b_{1}\leq\cdots\leq b_{m}$. Then $\sum_{k=1}^{\ell}a_{k}\geq\sum_{k=1}^{\ell}b_{k}$ and $\sum_{k=\ell}^{m}a_{k}\geq\sum_{k=\ell}^{m}b_{k}$ for all $\ell$. That is, $\{a_{1},\ldots,a_{m}\}$ (weakly) majorizes $\{b_{1},\ldots,b_{m}\}$ from both sides.
The majorization inequalities hold for $\ell=1$ by 3.9(ii) and $\ell=m$ by (3.7). These majorization inequalities have also been checked for multisets with $7$ elements and largest entry at most $15$. Out of the $\binom{15+7-1}{7}^{2}=13521038400$ pairs of multisets, $70653669$ or roughly $0.5\%$ yield cyclotomic generating functions. Of these, $2713$ do not satisfy $a_{k}\geq b_{k}$ for all $k$, or roughly $0.004\%$ of the cyclotomic generating functions. Each of these nonetheless satisfy the majorization inequalities in 3.12. Moreover, 3.12 holds for all $f$ of degree $\leq 42$, of which there are $10439036$. Finally, the majorization condition is preserved under multiplication.
Remark 3.13.
The affirmative answer to 3.12 together with Karamata’s inequality would give $\sum_{k=1}^{\ell}\psi(a_{k})\geq\sum_{k=1}^{\ell}\psi(b_{k})$ and $\sum_{k=\ell}^{m}\psi(a_{k})\geq\sum_{k=\ell}^{m}\psi(b_{k})$ for every non-decreasing convex function $\psi\colon[1,\infty)\to\mathbb{R}$, in particular strengthening (3.7) and leading to more necessary conditions for CGF’s.
3.5. CGF’s with a fixed denominator

### v3

@@ 3.2 Rational products of $q$,qintegers
$f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}=\sum_{k=0}^{\infty}c_{k}q^{k}$ (3.2)
as formal power series in $\mathbb{Z}[[q]]$. Such rational products include the set of basic CGFs. We examine several properties of such products.
In the next lemma, we state an explicit formula for the coefficients of the expansion of (3.2) generalizing work of Knuth for the number of permutations with $k\leqslant n$ inversions in $S_{n}$ in [Knu73, p.16]. See also [Sta12, Ex. 1.124] and [OEI23, A008302], and the application to standard Young tableaux in [BKS20b]. Here we use
@@
for all $k\in\mathbb{Z}_{\geqslant 0}$ and $x\in\mathbb{Z}$, including $x<0$. For all $x\in\mathbb{Z}$, define the empty product $\binom{x}{0}=1$.
Lemma 23.
@@
$c_{k}=\sum_{|\mu|=k}\prod_{\begin{subarray}{c}i\geqslant 1\\ m_{i}(\mu)>0\end{subarray}}\binom{M_{i}+m_{i}(\mu)-1}{m_{i}(\mu)}$ (3.3)
where $m_{i}(\mu)$ is the number of parts of $\mu$ of size $i$. Moreover, we may restrict the sum in (3.3) to only those $|\mu|=k$ where for all $i\geqslant 1$, either $M_{i}>0$ or $m_{i}(\mu)\leqslant|M_{i}|$.
@@
Equation (3.3) follows using the expansion $(1-q^{i})^{-j}=\sum_{n=0}^{\infty}\binom{j+n-1}{n}q^{in}$, multiplication of ordinary generating functions, and the fact that $\binom{x}{0}=1$. For the restriction, consider a term indexed by $\mu$ with $M_{i}\leqslant 0$ and $m_{i}(\mu)>|M_{i}|=-M_{i}$. It follows that $M_{i}\leqslant 0\leqslant M_{i}+m_{i}(\mu)-1$, so that
$m_{i}(\mu)!\binom{M_{i}+m_{i}(\mu)-1}{m_{i}(\mu)}=(M_{i}+m_{i}(\mu)-1)\cdots(M_{i}+2)(M_{i}+1)M_{i}=0.$
Remark 24.
The binomial coefficients $\binom{M_{i}+m_{i}(\mu)-1}{m_{i}(\mu)}$ appearing in (3.3) are positive when $M_{i}>0$. When $M_{i}\leqslant 0$ and $m_{i}(\mu)\leqslant|M_{i}|$, the binomial coefficient is nonzero and has sign $(-1)^{m_{i}(\mu)}$. Thus the negative summands in (3.3) are precisely those for which $\mu$ has an odd number of row lengths $i$ such that $M_{i}\leqslant 0$.
Remark 25.
In the particular case when $f(q)$ is a basic cyclotomic generating function, the fact that the expression in (3.3) is nonnegative, symmetric, and eventually 00 is quite remarkable. Is there a simplification of the expression in (3.3) that would directly imply nonnegativity of the coefficients for any $f(q)\in\Phi^{+}$? Even the case of $[n]_{q}!$ or any of the examples CGFs mentioned in the Introduction would be of interest.
3.3 Necessary conditions for $q$integer ratios to be polynomial
@@
$f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}$ (3.4)
to yield a polynomial (not necessarily with positive coefficients) is for the multiplicity of each primitive $d$th root of unity to be nonnegative. However, both numerator and denominator can be factored uniquely into cyclotomic polynomials, so the polynomiality criterion is equivalent to
#$\#\{k:\ell\mid a_{k}\}\geqslant\#\{k:\ell\mid b_{k}\}\hskip 7.22743pt\qquad\forall\ \ell\in\mathbb{Z}_{\geqslant 2}.$ (3.5)
Example 26.
Recall from the introduction that Stanley’s $q$-analogue of the hook length formula
@@
gives an important family of cyclotomic generating functions. In this case, the inequality (3.5) reduces to the fact that the $\ell$-quotient of any partition $\lambda$ with $|\lambda|=n$ has no more than $\lfloor n/\ell\rfloor$ cells. In particular, this gives a quick proof that $[n]_{q}!/\prod_{c\in\lambda}[h_{c}]_{q}\in\mathbb{Z}[q]$, though the stronger result that the coefficients are nonnegative is not clear from this approach. See [Sta79, Prop. 4.11] and [Sta99, Cor. 21.5] for more details and proofs for this equality. [BKS20a] for
3.4 Necessary conditions for a $q$integer ratio to be a CGF
In addition to satisfying the polynomiality conditions in (3.5) and (3.7), what further restrictions on $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{m}\}$ are required for a rational product of $q$-integers as in (3.4) to yield a cyclotomic generating function? Say $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/\prod_{k=1}^{m}[b_{k}]_{q}=\sum c_{k}q^{k}$. Using the expression for $c_{k}$ in (3.3), $c_{k}\geqslant 0$ and $\lim_{k\to\infty}c_{k}=0$ are by definition necessary and sufficient to prove $f(q)\in\Phi^{+}$, though these conditions are difficult to use for families of such rational products in practice. We will outline some more direct necessary conditions here. We begin with complete classifications for $m=1,2$.
Lemma 27.
@@ 3.4 Necessary conditions for a $q$integer ratio to be a CGF
  2. 2.
    If $f(q)$ is a power series with nonnegative coefficients, then $a_{1},a_{2}\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{b_{1},b_{2}\}$.
  3. 3.
@@
Conversely, suppose one of (i)-(iv) holds. Then, $f(q)$ is a polynomial if conditions (i) or (ii) hold by Lemma 27, and (iv) is equivalent to (iii), so it suffices to assume (iii) holds. Use the cyclotomic expansion of $[n]_{q}$ again from (2.3). Since (iii) holds, every cyclotomic divisor of either $[b_{1}]_{q}$ or $[b_{2}]_{q}$ but not both is canceled by a cyclotomic divisor of $[a_{1}]_{q}$, and every divisor of both $[b_{1}]_{q}$ and $[b_{2}]_{q}$ is canceled by divisors of $[a_{1}]_{q}$ and $[a_{2}]_{q}$ together. Thus, (1) holds.
The span condition in statement (2) is proved for all basic CGFs in Lemma 30 below. To prove (3), observe that $f(q)\in\Phi^{+}$ implies both the divisibility and span conditions hold by (1) and (2). Conversely, if any one of the divisibility conditions hold, then $f(q)$ is a polynomial with integer coefficients. If divisibility conditions (i) or (ii) hold, then $f(q)\in\Phi^{+}$ by Lemma 27, so suppose divisibility condition (iii) and the span condition in (2) holds. Again the case (iv) is similar.
By canceling out common cyclotomic factors, we may suppose without loss of generality that $\gcd(b_{1},b_{2})=1$ and that $a_{1}=b_{1}b_{2}$. Since $a_{2}\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{b_{1},b_{2}\}$, there exist positive integers $u,v$ such that $a_{2}=ub_{1}+vb_{2}$. So, we must show that the polynomial $f(q)=[b_{1}b_{2}]_{q}[ub_{1}+vb_{2}]_{q}/[b_{1}]_{q}[b_{2}]_{q}$ has nonnegative coefficients.
@@
Let $N(n)=\#\{x,y\in\mathbb{Z}_{\geqslant 0}:xb_{1}+yb_{2}=n\}$. The coefficient of $q^{n}$ in $f(q)$ is hence
@@
Let $N(n;\alpha,\beta)=\#\{\alpha\leqslant x,\beta\leqslant y:xb_{1}+yb_{2}=n\}$. The previous expression becomes
@@
Furthermore, let $N(n;\alpha,\beta;\delta)=\#\{\alpha\leqslant x<\alpha+\delta,\beta\leqslant y:xb_{1}+yb_{2}=n\}$. Thus, the expression for the coefficient of $q^{n}$ in $f(q)$ becomes
@@
Assume there exist $(x_{0},y_{0})\in\mathbb{Z}^{2}$ such that $x_{0}b_{1}+y_{0}b_{2}=n$, since otherwise the coefficient of $q^{n}$ in $f(q)$ is 00. Then, all integer solutions of $xb_{1}+yb_{2}=n$ are of the form $x=x_{0}-tb_{2},y=y_{0}+tb_{1}$ for some $t\in\mathbb{Z}$ by the theory of linear Diophantine equations. In particular, there exist unique solutions $(x_{1},y_{1}),(x_{2},y_{2})\in\mathbb{Z}^{2}$ with $0\leqslant x_{1}<b_{2}$ and $u\leqslant x_{2}<u+b_{2}$. Hence, $N(n;0,0;b_{2})=\delta_{y_{1}\geqslant 0}$ and $N(n;u,v;b_{2})=\delta_{y_{2}\geqslant v}$. Since $x_{1}\leqslant x_{2}$, we have $y_{1}\geqslant y_{2}$ as solutions to the linear equation. Thus, $\delta_{y_{2}\geqslant v}=1\Rightarrow\delta_{y_{1}\geqslant 0}=1$. Therefore, $N(n;0,0;b_{2})-N(n;u,v;b_{2})$ is nonnegative, which completes the proof. ∎
Example 29.
For $m=2$, the polynomiality condition alone does not imply nonnegative integer coefficients. For example, $[1]_{q}[6]_{q}/[2]_{q}[3]_{q}=\Phi_{6}(q)=q^{2}-q+1$. Here $6\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{2,3\}$, but $1$ is not. For $m=3$ and higher, the polynomiality and span conditions do not imply nonnegative integer coefficients. For example,
$\frac{[4]_{q}[4]_{q}[15]_{q}}{[2]_{q}[3]_{q}[5]_{q}}=q^{13}+q^{11}+q^{10}-q^{9}+2q^{8}+2q^{5}-q^{4}+q^{3}+q^{2}+1.$ (3.6)
However, we do have the following necessary conditions for basic CGFs related to spans and sums.
In the next lemma, we will use the big Theta notation for asymptoticly bounding a function. If $f(n)$ is a function of $n$, we write $f=\Theta(g)$ provided there exist constants $c,d$, a function $g(n)$, and a positive integer $n_{0}$ such that for all $n\geqslant n_{0}$ we have $cg(n)\leqslant f(n)\leqslant dg(n)$. Thus, the function $g$ is a reasonable approximation to $f$ for large $n$. We will extend the big Theta notation to other families of functions not necessarily indexed by positive integers if such constants exist.
Lemma 30.
If $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+}$, then the following hold.
@@
    For each $1\leqslant k\leqslant m$$a_{k}\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{b_{1},\ldots,b_{m}\}$.
@@
    If $a_{1}\leqslant\cdots\leqslant a_{m}$ and $b_{1}\leqslant\cdots\leqslant b_{m}$, then $a_{1}\geqslant b_{1}\geqslant 1$ and $a_{m}\geqslant b_{m}$.
@@
    Let $\mathcal{X}$ be the random variable corresponding to $f(q)/f(1)$ with mean $\mu=\frac{1}{2}\sum_{k=1}^{m}(a_{k}-b_{k})$ and variance $\sigma^{2}=\frac{1}{6}\sum_{k=1}^{m}(a_{k}^{2}-b_{k}^{2})$. Then, $\log\sigma=\Theta(\log\mu)$. More precisely,
    $\mu/2\leqslant\sigma^{2}\leqslant\mu^{2}.$
@@
    We have
    $\frac{\sum_{i=1}^{m}(a_{i}^{4}-b_{i}^{4})}{\left(\sum_{i=1}^{m}(a_{i}^{2}-b_{i}^{2})\right)^{2}}\leqslant\frac{5}{3}.$
@@
    $\sum_{k=1}^{m}a_{k}^{d}\geqslant\sum_{k=1}^{m}b_{k}^{d},$ (3.7)
    where the inequality is strict if the polynomial is non-constant. For example, when $d=1$ this inequality is equivalent to noting the degree of $f(q)$ is $\sum_{k=1}^{m}a_{k}-b_{k}\geqslant 0$.
@@
The support of the power series $\prod_{k=1}^{m}1/(1-q^{b_{k}})$ is $\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{b_{1},\ldots,b_{m}\}$. Since $f(q)$ has nonnegative coefficients and constant term $1$, comparing supports gives
$\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{a_{1},\ldots,a_{m}\}\subset\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{b_{1},\ldots,b_{m}\},$
which yields (i).
(ii) From (i), it follows that $a_{1}\geqslant\min\{b_{1},\ldots,b_{m}\}=b_{1}$. From (3.5) at $\ell=b_{m}$, there is some $a_{k}$ such that $b_{m}\mid a_{k}\leqslant a_{m}$, so $b_{m}\leqslant a_{m}$.
@@
(iv) By definition, the kurtosis of $\mathcal{X}$ is $\mathbb{E}[(\frac{X-\mu}{\sigma})^{4}]$. Jensen’s Inequality for convex functions gives the inequality of central moments
$\alpha_{2}^{2}=\mathbb{E}[(\mathcal{X}-\mu)^{2}]^{2}\leqslant\mathbb{E}[(\mathcal{X}-\mu)^{4}]=\alpha_{4}.$
In terms of cumulants, this says $\kappa_{2}^{2}\leqslant\kappa_{4}+3\kappa_{2}^{2}$, which simplifies to the stated expression by (2.5).
(v) The inequality follows directly from the formula for cumulants of a rational product of $q$-integers (2.5) and the sign constraints in Corollary 17. ∎
Remark 31.
@@
Remark 32.
One might wonder if $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+}$ implies $a_{k}\geqslant b_{k}$ for all $k$ when sorted as in Lemma 30(ii). However, this is false, though empirically counterexamples are rare. For instance,
@@
is one of only three counterexamples with $5$ $q$-integers in the numerator and denominator and maximum entry $\leqslant 14$, out of $956719$ cyclotomic generating functions satisfying these conditions. On the other hand, we propose the following.
Conjecture 33.
Suppose $f(q)=\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+}$$a_{1}\leqslant\cdots\leqslant a_{m}$,  $b_{1}\leqslant\cdots\leqslant b_{m}$. Then $\sum_{k=1}^{\ell}a_{k}\geqslant\sum_{k=1}^{\ell}b_{k}$ and $\sum_{k=\ell}^{m}a_{k}\geqslant\sum_{k=\ell}^{m}b_{k}$ for all $\ell$. That is, $\{a_{1},\ldots,a_{m}\}$ (weakly) majorizes $\{b_{1},\ldots,b_{m}\}$ from both sides.
The majorization inequalities hold for $\ell=1$ by Lemma 30(ii) and $\ell=m$ by (3.7). These majorization inequalities have also been checked for multisets with $7$ elements and largest entry at most $15$. Out of the $\binom{15+7-1}{7}^{2}=13521038400$ pairs of multisets, $70653669$ or roughly $0.5\%$ yield cyclotomic generating functions. Of these, $2713$ do not satisfy $a_{k}\geqslant b_{k}$ for all $k$, or roughly $0.004\%$ of the cyclotomic generating functions. Each of these nonetheless satisfy the majorization inequalities in 33. Moreover, 33 holds for all $f$ of degree $\leqslant 42$, of which there are $10439036$. Finally, the majorization condition is preserved under multiplication.
Remark 34.
The affirmative answer to 33 together with Karamata’s inequality would give $\sum_{k=1}^{\ell}\psi(a_{k})\geqslant\sum_{k=1}^{\ell}\psi(b_{k})$ and $\sum_{k=\ell}^{m}\psi(a_{k})\geqslant\sum_{k=\ell}^{m}\psi(b_{k})$ for every non-decreasing convex function $\psi\colon[1,\infty)\to\mathbb{R}$, in particular strengthening (3.7) and leading to more necessary conditions for CGFs.
3.5 CGFs with a fixed denominator